QUESTION NINE


Malawi and Zambia both produce masks and sanitisers. Suppose that a Malawian worker can produce


50 masks per hour or 1 sanitiser per hour. Suppose that a Zambian worker can produce 40 masks per


hour or 2 sanitisers per hour.


a. Which country has the absolute advantage in the production of each good?


b. Which country has the comparative advantage?


c. If Malawi and Zambia decide to trade, which commodity will Zambia trade to Malawi? Explain.


d. If a Zambian worker could produce only 1 sanitiser per hour, would Zambia still gain from trade?


Would Malawi still gain from trade? Explain.[15 marks]

1
Expert's answer
2022-01-03T09:23:58-0500

a. Malawi has an absolute advantage in the production of masks, and Zambia in the production of disinfectants

b.

Malawi

"50masks=\frac{1}{50}sanitiser"

Zambia 

"40masks=\frac{2}{40}sanitiser"

"0.02sanitiser<0.05sanitiser"

Malawi has a comparative advantage in masks, and Zambia accordingly has a comparative advantage in the production of disinfectants.

с. Consequently, Malawi will export surplus masks and import sanitizers, and Zambia will export sanitizers and import masks.

d.Malawi

"50masks=\frac{1}{50}sanitiser"

Zambia 

"40masks=\frac{1}{40}sanitiser"

"0.02sanitiser<0.025sanitiser"

Malawi has a comparative advantage in masks, and Zambia accordingly has a comparative advantage in the production of disinfectants.

Consequently, Malawi will export surplus masks and import sanitizers, and Zambia will export sanitizers and import masks.

Malawi will only benefit from the mask trade